OpenStudy (anonymous):

How do you solve: 5x^3/10x^7 Steps please :) Thank you.

OpenStudy (anonymous):

\[5x^3/10x^7\]

OpenStudy (alexwee123):



OpenStudy (unklerhaukus):

\[\frac{5x^3}{10x^7=\frac{5}{10}x^{3-7}\]

OpenStudy (alexwee123):